# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v1

@@ Abstract
Model compression is essential in the deployment of large Computer Vision models on embedded devices. However, static optimization techniques (e.g. pruning, quantization, etc.) neglect the fact that different inputs have different complexities, thus requiring different amount of computations. Dynamic Neural Networks allow to condition the number of computations to the specific input. The current literature on the topic is very extensive and fragmented. We present a comprehensive survey that synthesizes and unifies existing Dynamic Neural Networks research in the context of Computer Vision. Additionally, we provide a logical taxonomy based on which component of the network is adaptive: the output, the computation graph or the input. Furthermore, we argue that Dynamic Neural Networks are particularly beneficial in the context of Sensor Fusion for better adaptivity, noise reduction and information prioritization. We present preliminary works in this direction.
@@ 1 Introduction
In the past decade, the introduction of architectural paradigms such as Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) has enabled substantial advancements in the Computer Vision field. The use of these models for practical applications has also driven the necessity to explore reduction techniques for deployment in low-computation devices. Model compression methods like quantization (Gholami et al., 2021), pruning (Cheng et al., 2024), knowledge distillation (Gou et al., 2021), neural architecture search (White et al., 2023), and low-rank factorization (Swaminathan et al., 2020) have demonstrated how Neural Networks are often unnecessarily overparameterized  (Deng et al., 2020). On the other hand, these optimization techniques do not take into account the different amount of computations needed for each input, disregarding the fact that different inputs have different complexities. Dynamic Neural Networks allow one to work around this problem by conditioning the computation path of a network to its input—refer to Figure 1 for an overview of the different types of Dynamic Neural Networks.
@@
A parallel can be drawn with the human perception system, which is able to adaptively allocate time and scrutiny for visual recognition, where simple glimpses are sufficient to recognize simple objects and scenes, while more time and attention are required in cases where the elements are occluded or complicated (Schiebener et al., 2013, Walther et al., 2010).
A Dynamic Neural Network is defined as a deep learning model that is able to adjust the number of computations according to the complexity of the input (Han et al., 2021, Xu and McAuley, 2023). This is obtained by introducing some elasticity within the network (usually gating or decision modules), which allows one to differentiate between easy and hard samples and take a computation path decision accordingly. In some cases, the complexity estimation can also be steered by other external factors such as the overall computational budget or the environmental conditions (Wang et al., 2019, Malawade et al., 2022, Devoto et al., 2024).
@@
Figure 2: Overview of the publications considered in this survey, grouped by year and topic. In total, 148 publications have been reviewed: 62 for the Early Exits Section, 44 for the Dynamic Routing Section, 27 for the Token Skimming Section, and 15 in the Dynamic Sensor Fusion Section.
Thus, we present a survey of Dynamic Neural Networks in the context of Computer Vision, outlining the most representative methods. We categorize concepts, methods, and techniques in this field through the structuring of clear taxonomy so that it becomes easier to understand and navigate. Our goal is to create a standardized language that supports the exploration of methodologies and applications in the domain through a comparison of recent advancements.
@@
forked edges, folder indent=1cm, where=level()¡1folder, grow’=east, align=center, where=level()¿0l sep+=1.5cm, for tree= fork sep=1.5mm, thick, edge=thick, font=, align=left, if n children=0 if level=2yshift=-5mm, for parent=s sep=1.5mm draw , base1 [A Survey on Dynamic Neural Networks:
from Computer Vision to
Multi-modal Sensor Fusion, calign=edge midpoint, s sep=2cm, font=, fill=white, align=center, draw [2 Dynamic Computer Vision Techniques, fill=greyset [2.1 Early Exits, fill=greenset [2.1.1 Architectures, base1, fill=greenset!75] [2.1.2 Methods, draw=none, base1, fill=greenset!75, [Exit policy, fill=greenset!35] [Loss weighting, fill=greenset!35] [Exiting positioning, fill=greenset!35] [Learning methods, fill=greenset!35] [Train-test mismatch, fill=greenset!35] [Self distillation, fill=greenset!35] [Bag-of-Features, quantization, OOD, and NAS, fill=greenset!35] ] [2.1.3 Applications, base1, fill=greenset!75] ] [2.2 Computational Routing, fill=blueset [2.2.1 Dynamic architectures, base1, draw=none, fill=blueset!75 [Dynamic computation path, fill=blueset!35] [Dynamic components, fill=blueset!35] ] [2.2.2 Mixture of Experts, base1, fill=blueset!75] [2.2.3 Applications, base1, fill=blueset!75] ] [2.3 Token Skimming, fill=redset [2.3.1 Token dropping, base1, draw=none, fill=redset!75, [Hierarchical, fill=redset!35] [Per block, fill=redset!35] ] [2.3.2 Token merging, base1, fill=redset!75] [2.3.3 Other methods and applications, base1, fill=redset!75] ] ] [3 Applications in Sensor Fusion, fill=yellowset, [Object detection, fill=yellowset!35] [Semantic segmentation, fill=yellowset!35] [Autonomous vision-based navigation, fill=yellowset!35] [Other sensor fusion
approaches, fill=yellowset!35] ] ]
@@
This survey discusses Dynamic Neural Networks only in the context of Computer Vision and their application in Sensor Fusion, independently of the task that the paper tackles. We do not cover papers coming from other fields in deep learning (e.g. Natural Language Processing, Graph Neural Networks), although some of those might be mentioned when important in the context of the discussion. Furthermore, we considered only papers from the last 8 years, i.e., from January 2016 to September 2024.
@@
This paper is related to existing surveys: Scardapane et al. (2020) discusses the theoretical foundation theory for Early Exits Neural Networks. Chen et al. (2021a) explores sparsity in ViT models, including Token Skimming approaches discussed in detail also in our survey, while Farina et al. (2024) is a more recent survey on the same topic. Han et al. (2021) discusses advancements in Dynamic Routing techniques, focusing mostly on the dynamicity of the computational path. Scardapane et al. (2024) presents a general formalization of all these approaches, which tries to unify under the same definition all conditional computing methods. Matsubara et al. (2022) provides a survey that puts in relation Early Exits models with computation splitting between on-device and cloud resources. In Xu and McAuley (2023), a survey on Dynamic Neural Networks for Natural Language Processing is presented. This paper shares a very similar structure with ours but focuses on NLP and presents a limited selection of publications.
Finally, to the best of our knowledge, no existing survey comprehensively examines the application of Dynamic Neural Networks in the context of Sensor Fusion.
@@
In contrast to existing works, our survey paper provides a comprehensive overview of relevant papers in the domain of Dynamic Neural Networks for Computer Vision, while at the same time, linking them to the highly relevant field of Sensor Fusion. We can summarize our main contributions as follows:
  • We present a comprehensive and up-to-date survey focusing on Computer Vision techniques that adjust the number of computations of their architecture dynamically, based on the input complexity. We present a total of 146 publications relevant to the field.
  • We provide a clear taxonomy of the methods according to which component of the network is adaptable. We grouped the sections according to where the network provides output (Early Exits), how the network builds its computation path (Dynamic Routing), or what portion of the input has been considered (Token Skimming).
  • We argue that Dynamic Neural Network techniques are especially beneficial in the context of Sensor Fusion. We present all the seminal works available in this regard, dividing them according to the task they tackle.
@@ 2.1 Early exits
The idea at the base of the development of Early Exits networks is rooted in the assumption that samples in a dataset can vary a lot in complexity. Following this principle, traditional Neural Networks tend to be overparametrized for most of the samples, which could potentially be inferred with just a fraction of the parameters used. Furthermore, in the case of easy samples, Convolutional Neural Networks (CNNs) have been shown to be capable of extracting features in the early layers that are already sufficient for a confident classification (Panda et al., 2016). For these reasons, Early Exits networks propose to evaluate the confidence of the model in the prediction of output at different stages within the network, with the option of exiting the computation path earlier, if the model is confident enough on the prediction. This has a twofold purpose: it allows to optimize the computational cost of the model tailored to the input difficulty, and moreover, it avoids the phenomenon of overthinking, a circumstance in which a correct prediction in the early stages of a network gets corrupted by noise in deeper layers (Kaya et al., 2019). The confidence evaluation of a model is carried out by attaching a number of auxiliary classifiers to the intermediate blocks of the network. Section 2.1.2 further illustrates different methods for the determination of the number and position of the exit points, and how to properly construct and evaluate the prediction confidence score of a network in the early stages. Before exploring the major contributions in the field, we present here a summary of the formulation of the Early Exits models presented in Scardapane et al. (2020) and Scardapane et al. (2024), to which we refer for more details.
@@ A Neural Network can be seen as a composition of $b$ sequential blocks:
We proceed now to illustrate the major contributions in the field of Early Exits models. Section 2.1.1 presents the key architectures that have been introduced in the literature, Section 2.1.2 discusses papers that try to improve specific aspects of the learning process of these types of networks, while Section 2.1.3 focuses on applications to other tasks rather than classification.
@@ 2.1.1 Architectures
Yang et al. (2020) proposes some variation in the implementation with respect to MSDNet. Compared to MSDNet, the dense blocks with the smallest scale input are taken into account first and processed throughout all depths of the network. The processing at finer scales happens at a later stage. The key idea is that this low-scale resolution is sufficient to classify easy samples better in the early stages, whereas higher resolution is beneficial only for more difficult inputs later on, which are also more computationally costly.
More recently, with the advent of Transformer-based architectures, Wang et al. (2021) proposes a framework named Dynamic Vision Transformer (DVT) which aims at finding the correct patch size for tokenization at which the image has to be processed to guarantee a confident output. In the Transformer architecture, the complexity of the model is quadratic with respect to the number of tokens that attend the attention block. For this reason, a lower number of tokens translates into a much faster model. The paper trains a cascade of Transformers, where each model uses a larger number of smaller tokens at each scale. In this way, images are processed at a coarse level at first with a low number of tokens, and then more details are processed at each iteration including a larger number of tokens derived from smaller patches. The model exits in the case where the prediction output is sufficiently confident. The paper presents also feature-wise and relationship-wise reuse mechanisms to reduce redundant computations already performed by the layer that processed the coarser tokens.
@@ 2.1.2 Methods
The exit policy is the logical component within an Early Exits network which is in charge of determining whether the model has processed the input enough to provide an intermediate output. Scardapane et al. (2020) proposes a method to optimize the Early Exits decision jointly during training by adding a soft gate at each exit that decides a weighting strategy. The idea is to learn a weighting scheme to combine the output of the current prediction with the (supposedly improved) output of the next exit’s prediction. Essentially, the key idea is for the network to self-weight how much it should rely on the output of the current exit with respect to the output of the next Early Exit. This strategy allows training the model by exploiting a cross-entropy loss with respect to the final output solely; thus, it does not need to create a loss that is a weighted combination of all the losses computed on the intermediate exits. During inference, the added auxiliary gating value can be considered a standard binary classifier and can be thresholded to evaluate the probability of exit. Furthermore, the paper proposes a regularization term for the final loss that induces the training process to take into account also the average computational cost of the network. Similarly, in Chen et al. (2020) the same idea of learning the stopping policy for Early Exits together with the network parameters is considered. In this case, the solution views the stopping time as a latent variable conditioned on the input, interpreting the problem as a variational Bayes one. The training procedure to achieve this consists of decomposing the problem in two stages, necessary for learning both the predictive model and the stopping policy. In the first stage the model parameters are learned, assuming that the stop time always follows the best stopping distribution. In the second stage, the exit policy is learned by minimizing the reverse Kullback–Leibler divergence through a Reinforcement Learning algorithm.
@@
Loss weighting methods try to determine the correct contribution of each exit output to the overall loss, with the aim of improving the learning of the network. It consists in determining the $\alpha_{i}$ values in equation (3). Wang et al. (2019) proposes DynExit, a dynamic early-exit strategy for ResNet (He et al., 2015). This strategy introduces a dynamic loss-weight adjustment mechanism contributing to the weighting of the final loss as a weighted sum of all the losses, each computed at the respective exit. Instead of using manually tuned weighting values, in DynExit the weights become a trainable parameter with a dynamic range, which enables the loss-weight of different exit points. For this reason, the correct loss combination can be learned with all the other parameters, thus pondering the overall contribution of single exits to the learning process and increasing the overall accuracy. Furthermore, some hardware-aware optimizations of the implementation of the softmax and the cross entropy loss are presented, which are adapted to achieve high speed with low hardware complexity.
@@
In this paragraph, we cover those techniques that focus on the determination of the optimal amount and position of the exit points within an Early Exits network. The positioning of the exits is a straightforward problem where in most cases the default option is to have the exits equally distributed within the depth of the network. Nonetheless, some studies have researched the optimal position of these exit points. Laskaridis et al. (2020) presents HAPI, a methodology for creating an optimal Early Exits network design by optimizing the placement of the intermediate exits as well as the number of exits within a network. The Early Exits variants of the network are represented in a graph-matrix form that allows the design to be solved as a mathematical optimization problem. The method changes the number and position of Early Exits along the network by adding and removing nodes in the network graph.
Recently, Liu et al. (2024) proposes a solution to add Early Exits to ViT models, called MET. It introduces mainly two novelties: first, in between Transformer blocks, a devised adapter extracts appropriate representations for different exits in a shared space. To reduce the number of trainable parameters, different adapters in the network share a part of the parameters. Furthermore, supervised graphs, which characterize the intra-class and inter-class relationships among samples, are adopted to regularize the representation-learning processes of Early Exits. Furthermore, Li et al. (2023) builds on top of the work of Passalis et al. (2020) by allowing a fine-grained choice on the exit points; instead of fixing a determined amount of exit points, a learned component decides where to place optimally the position of classifiers for Early Exiting with respect to any block in the network. This approach takes also into account the trade-off between model accuracy and energy cost through the use of a hyperparameter. If the classification determines that the network needs more computation to get to a certain result, the next exit point position will be predicted. Finally, Lin et al. (2024) also further analyzes the effect of different design choices on the positioning of the branches.
@@
Finally, the Zero Time Waste (ZTW) (Wołczyk et al., 2021, Wójcik et al., 2023) approach proposes to reuse the output of internal classifiers as inductive prediction information to the later classification blocks of the model, for the cases where the exiting has been contemplated by performing a classification, but not performed. This is done by adding direct connections between classifiers and by combining previous outputs in an ensemble-like manner. Part of the work also demonstrates how the method performance can be improved with knowledge distillation between internal classifiers.
@@
More recently, also Regol et al. (2024) arrives at the same solution with a slightly different technique. The authors present a learning method for the gating and inference mechanisms. The paper proposes to optimize a loss that jointly assesses accuracy and inference cost by formulating the minimization as a bilevel optimization task. Each level of the bilevel optimization is simpler than the overall problem, leading to more stable learning.
@@
Finally, Zhang et al. (2024) also proposes to train through self-distillation a multi-exit network. The key contributions of this paper are a dynamic generation of the importance weights for each classifier, a logit-based self-distillation module, and a feature-based self-distillation module.
@@
Quantization is the process of reducing the precision of weights and activations from higher bit-widths to lower ones, generally to reduce the model size and improve inference speed. Regol et al. (2024) proposes QuEE, a solution that combines quantization and Early Exits dynamic network. Rather than gating between exiting or continuing the computations, the decision verges among continuing with full or reduced computation. The proposed architecture allows to control both the number of blocks that are evaluated, through the use of Early Exits, and the amount of computation conducted within each block. This is conducted by using a clustering algorithm over a discretized space of possibilities.

### v3

@@ Abstract
Model compression is essential in the deployment of large Computer Vision models on embedded devices. However, static optimization techniques (e.g. pruning, quantization, etc.) neglect the fact that different inputs have different complexities, thus requiring different amount of computations. Dynamic Neural Networks allow to condition the number of computations to the specific input. The current literature on the topic is very extensive and fragmented. We present a comprehensive survey that synthesizes and unifies existing Dynamic Neural Networks research in the context of Computer Vision. Additionally, we provide a logical taxonomy based on which component of the network is adaptive: the output, the computation graph or the input. Furthermore, we argue that Dynamic Neural Networks are particularly beneficial in the context of Sensor Fusion for better adaptivity, noise reduction and information prioritization. We present preliminary works in this direction. We complement this survey with a curated repository listing all the surveyed papers, each with a brief summary of the solution and the code base when available: https://github.com/DTU-PAS/awesome-dynn-for-cv.
@@ 1 Introduction
In the past decade, the introduction of architectural paradigms such as Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) has enabled substantial advancements in the Computer Vision field. Efficiency then becomes an important factor for deploying these neural networks—especially in the foundation model era. Model compression methods like quantization (Gholami et al., 2021), pruning (Cheng et al., 2024), knowledge distillation (Gou et al., 2021), neural architecture search (White et al., 2023), and low-rank factorization (Swaminathan et al., 2020) have demonstrated how Neural Networks are often unnecessarily overparameterized  (Deng et al., 2020). On the other hand, these optimization techniques do not take into account the different amount of computations needed for each input, disregarding the fact that different inputs have different complexities. Dynamic Neural Networks allow one to work around this problem by conditioning the computation path of a network to its input—refer to Figure 1 for an overview of three different types of Dynamic Neural Networks which are the focus of this survey paper.
@@
A parallel can be drawn with the human perception system, which is able to adaptively allocate time and scrutiny for visual recognition, where simple glimpses are sufficient to recognize simple objects and scenes, while more time and attention are required in cases where the elements are occluded or complicated (Schiebener et al., 2013; Walther et al., 2010).
A Dynamic Neural Network is defined as a deep learning model that is able to adjust the number of computations according to the complexity of the input (Han et al., 2021; Xu and McAuley, 2023). This is obtained by introducing some elasticity within the network (usually gating or decision modules), which allows one to differentiate between easy and hard samples and take a computation path decision accordingly. In some cases, the complexity estimation can also be steered by other external factors such as the overall computational budget or the environmental conditions (Wang et al., 2019b; Malawade et al., 2022b; Devoto et al., 2024).
@@
Figure 2: Overview of the publications considered in this survey, grouped by year and topic. In total, 163 publications have been reviewed: 62 for the Early Exits Section, 45 for the Dynamic Routing Section, 27 for the Token Skimming Section, and 16 in the Dynamic Sensor Fusion Section.
Thus, we present a survey of Dynamic Neural Networks in the context of Computer Vision, outlining the most representative methods. We categorize concepts, methods, and techniques in this field through the structuring of a clear taxonomy so that it becomes easier to understand and navigate the field. Our goal is to create a standardized language that supports the exploration of methodologies and applications in the domain through a comparison of recent advancements.
@@
This survey discusses Dynamic Neural Networks only in the context of Computer Vision and their application in Sensor Fusion, independently of the task that the paper tackles. We do not cover papers coming from other fields in deep learning (e.g. Natural Language Processing, Graph Neural Networks), although some of those might be mentioned when important in the context of the discussion. Furthermore, we considered only papers from the last 9 years, i.e., from January 2016 to September 2025.
@@
This paper is related to existing surveys: Scardapane et al. (2020b) is the first survey in the field to set the base for a theoretical description of Early Exits Neural Networks, unifying different methodologies under a common formulation. Scardapane et al. (2024) updates and generalizes this formalization also to other Dynamic Neural Network approaches, i.e. Mixture-of-Experts and Token sparsification. In contrast, our survey takes a complementary approach where we focus on analyzing all relevant papers on the topic, rather than a theoretical unification of the field. The goal is to provide a more complete insight into all the available ideas and applications of these methods. Chen et al. (2021a) explores sparsity in ViT models, focusing on Token Skimming approaches, while Farina et al. (2024) is a more recent survey on the same topic. Han et al. (2021) discusses advancements in Dynamic Routing techniques, focusing mostly on the dynamicity of the computational path. All these survey papers highlight the most relevant works in both fields of Natural Language Processing and Computer Vision. On the other hand, in this survey, we deliberately consider only the Computer Vision field, in order to provide a more in-depth overview of all relevant publications and not only major milestones. Matsubara et al. (2022) provides a survey that puts in relation Early Exits models with computation splitting between on-device and cloud resources. In Xu and McAuley (2023), a survey on Dynamic Neural Networks for Natural Language Processing is presented. This paper shares a similar structure with ours but focuses on NLP and presents a limited selection of publications. We provide a survey that is complementary but focuses on Computer Vision papers.
Finally, to the best of our knowledge, no existing survey comprehensively examines the application of Dynamic Neural Networks in the context of Sensor Fusion. Since we consider this a relevant and emerging topic, we dedicate a chapter specifically to all the relevant publications in this subtopic.
@@
Unlike existing surveys, our work offers a comprehensive overview of Dynamic Neural Networks for Computer Vision, while simultaneously connecting them to the crucial field of Sensor Fusion. Our primary contributions are summarized as follows:
  • 1.
    We carry out a comprehensive and up-to-date survey of Computer Vision methods that dynamically adjust their computational load based on input complexity. In total, we review 163 publications relevant to the domain.
  • 2.
    We propose a clear taxonomy categorizing methods based on the adaptive network component. Specifically, we group approaches by where the network yields output (Early Exits), how it constructs its computation path (Dynamic Routing), or which portion of the input is processed (Token Skimming).
  • 3.
    We highlight that Dynamic Neural Network strategies are particularly advantageous for Sensor Fusion. We discuss the seminal works in this area, organizing them according to the specific task they address.
  • 4.
    Finally, we provide a selected online repository listing all referenced papers, accompanied by brief descriptions and links to the corresponding codebases (where available) for easy reference. The repository is available at: https://github.com/DTU-PAS/awesome-dynn-for-cv.
@@ 2.1 Early exits
The idea at the base of the development of Early Exits networks is rooted in the assumption that samples in a dataset can vary a lot in complexity. Following this principle, traditional Neural Networks tend to be overparametrized for most of the samples, which could potentially be inferred with just a fraction of the parameters used. Furthermore, in the case of easy samples, Convolutional Neural Networks (CNNs) have been shown to be capable of extracting features in the early layers that are already sufficient for a confident classification (Panda et al., 2016). For these reasons, Early Exits networks propose to evaluate the confidence of the model in the prediction of output at different stages within the network, with the option of exiting the computation path earlier, if the model is confident enough on the prediction. This has a twofold purpose: it allows to optimize the computational cost of the model tailored to the input difficulty, and moreover, it avoids the phenomenon of overthinking, a circumstance in which a correct prediction in the early stages of a network gets corrupted by noise in deeper layers (Kaya et al., 2019). The confidence evaluation of a model is carried out by attaching a number of auxiliary classifiers to the intermediate blocks of the network. Section 2.1.2 further illustrates different methods for the determination of the number and position of the exit points, and how to properly construct and evaluate the prediction confidence score of a network in the early stages. Before exploring the major contributions in the field, we present here a summary of the formulation of the Early Exits models presented in Scardapane et al. (2020b) and Scardapane et al. (2024), to which we refer for more details.
@@ A Neural Network can be seen as a composition of $b$ sequential blocks:
We proceed now to illustrate the major contributions in the field of Early Exits models. Section 2.1.1 presents the key architectures that have been introduced in the literature, Section 2.1.2 discusses papers that try to improve specific aspects of the learning process of these types of networks, while Section 2.1.3 focuses on applications to other tasks rather than classification. It should be noted that numerical results for the presented publications are not provided in this section. The domain of Early Exits lacks a standardized experimental protocol. We report on studies whose findings are based on a diverse array of datasets and evaluation architectures. A direct and fair quantitative analysis infeasible without extensive re-implementation. This heterogeneity highlights a critical gap in the current literature: the necessity for a structured benchmark. We argue that such a benchmark is essential to consistently compare different Dynamic Neural Network models, specifically accounting for inference speed and efficiency, the variability of computations in relation to input complexity, and performance quality in edge cases, such as small objects or targets located in the peripheral regions of the image. A further discussion on the topic is presented in the Discussion (Section 4).
@@ 2.1.1 Architectures
Yang et al. (2020) proposes some variation in the implementation with respect to MSDNet. Compared to MSDNet, the dense blocks with the smallest scale input are taken into account first and processed throughout all depths of the network. The processing at finer scales happens at a later stage. The key idea is that this low-scale resolution is sufficient to classify easy samples better in the early stages, whereas higher resolution is beneficial only for more difficult inputs later on, which are also more computationally costly. Also, Addad et al. (2025) further explores the use of MSDNet architecture for early exits by improving the feature representation and aggregation based on self-attention.
More recently, with the advent of Transformer-based architectures, Wang et al. (2021b) proposes a framework named Dynamic Vision Transformer (DVT) which aims at finding the correct patch size for tokenization at which the image has to be processed to guarantee a confident output. In the Transformer architecture, the complexity of the model is quadratic with respect to the number of tokens that attend the attention block. For this reason, a lower number of tokens translates into a much faster model. The paper trains a cascade of Transformers, where each model uses a larger number of smaller tokens at each scale. In this way, images are processed at a coarse level at first with a low number of tokens, and then more details are processed at each iteration including a larger number of tokens derived from smaller patches. The model exits in the case where the prediction output is sufficiently confident. The paper presents also feature-wise and relationship-wise reuse mechanisms to reduce redundant computations already performed by the layer that processed the coarser tokens.
@@ 2.1.2 Methods
The exit policy is the logical component within an Early Exits network which is in charge of determining whether the model has processed the input enough to provide an intermediate output. Scardapane et al. (2020a) proposes a method to optimize the Early Exits decision jointly during training by adding a soft gate at each exit that decides a weighting strategy. The idea is to learn a weighting scheme to combine the output of the current prediction with the (supposedly improved) output of the next exit’s prediction. Essentially, the key idea is for the network to self-weight how much it should rely on the output of the current exit with respect to the output of the next Early Exit. This strategy allows training the model by exploiting a cross-entropy loss with respect to the final output solely; thus, it does not need to create a loss that is a weighted combination of all the losses computed on the intermediate exits. During inference, the added auxiliary gating value can be considered a standard binary classifier and can be thresholded to evaluate the probability of exit. Furthermore, the paper proposes a regularization term for the final loss that induces the training process to take into account also the average computational cost of the network. Similarly, in Chen et al. (2020) the same idea of learning the stopping policy for Early Exits together with the network parameters is considered. In this case, the solution views the stopping time as a latent variable conditioned on the input, interpreting the problem as a variational Bayes one. The training procedure to achieve this consists of decomposing the problem in two stages, necessary for learning both the predictive model and the stopping policy. In the first stage the model parameters are learned, assuming that the stop time always follows the best stopping distribution. In the second stage, the exit policy is learned by minimizing the reverse Kullback–Leibler divergence through a Reinforcement Learning algorithm.
@@
Very recently, Bajpai and Hanawal (2025) proposes the previous exit responses as experts for an agreement that provides robustness in the exit confidence decision. If the previous exits were leaning toward a different class, the current expert is considered not confident, and the network continues the computations.
@@
Loss weighting methods try to determine the correct contribution of each exit output to the overall loss, with the aim of improving the learning of the network. It consists in determining the $\alpha_{i}$ values in equation (3). Wang et al. (2019a) proposes DynExit, a dynamic early-exit strategy for ResNet (He et al., 2016). This strategy introduces a dynamic loss-weight adjustment mechanism contributing to the weighting of the final loss as a weighted sum of all the losses, each computed at the respective exit. Instead of using manually tuned weighting values, in DynExit the weights become a trainable parameter with a dynamic range, which enables the loss-weight of different exit points. For this reason, the correct loss combination can be learned with all the other parameters, thus pondering the overall contribution of single exits to the learning process and increasing the overall accuracy. Furthermore, some hardware-aware optimizations of the implementation of the softmax and the cross entropy loss are presented, which are adapted to achieve high speed with low hardware complexity.
@@
In this paragraph, we cover those techniques that focus on the determination of the optimal amount and position of the exit points within an Early Exits network. The positioning of the exits is a straightforward problem where in most cases the default option is to have the exits equally distributed within the depth of the network. Nonetheless, some studies have researched the optimal position of these exit points. Laskaridis et al. (2020b) presents HAPI, a methodology for creating an optimal Early Exits network design by optimizing the placement of the intermediate exits as well as the number of exits within a network. The Early Exits variants of the network are represented in a graph-matrix form that allows the design to be solved as a mathematical optimization problem. The method changes the number and position of Early Exits along the network by adding and removing nodes in the network graph.
Recently, Liu et al. (2024c) proposes a solution to add Early Exits to ViT models, called MET. It introduces mainly two novelties: first, in between Transformer blocks, a devised adapter extracts appropriate representations for different exits in a shared space. To reduce the number of trainable parameters, different adapters in the network share a part of the parameters. Furthermore, supervised graphs, which characterize the intra-class and inter-class relationships among samples, are adopted to regularize the representation-learning processes of Early Exits. Furthermore, Li et al. (2023a) builds on top of the work of Passalis et al. (2020) by allowing a fine-grained choice on the exit points; instead of fixing a determined amount of exit points, a learned component decides where to place optimally the position of classifiers for Early Exiting with respect to any block in the network. This approach takes also into account the trade-off between model accuracy and energy cost through the use of a hyperparameter. If the classification determines that the network needs more computation to get to a certain result, the next exit point position will be predicted. Finally, Lin et al. (2024) also further analyzes the effect of different design choices on the positioning of the branches.
@@
Finally, the Zero Time Waste (ZTW) (Wołczyk et al., 2021; Wójcik et al., 2023) approach proposes to reuse the output of internal classifiers as inductive prediction information to the later classification blocks of the model, for the cases where the exiting has been contemplated by performing a classification, but not performed. This is done by adding direct connections between classifiers and by combining previous outputs in an ensemble-like manner. Part of the work also demonstrates how the method performance can be improved with knowledge distillation between internal classifiers.
@@
More recently, also Regol et al. (2024b) arrives at the same solution with a slightly different technique. The authors present a learning method for the gating and inference mechanisms. The paper proposes to optimize a loss that jointly assesses accuracy and inference cost by formulating the minimization as a bilevel optimization task. Each level of the bilevel optimization is simpler than the overall problem, leading to more stable learning.
@@
Finally, Zhang et al. (2024a) also proposes to train through self-distillation a multi-exit network. The key contributions of this paper are a dynamic generation of the importance weights for each classifier, a logit-based self-distillation module, and a feature-based self-distillation module.
@@
Quantization is the process of reducing the precision of weights and activations from higher bit-widths to lower ones, generally to reduce the model size and improve inference speed. Regol et al. (2024a) proposes QuEE, a solution that combines quantization and Early Exits dynamic network. Rather than gating between exiting or continuing the computations, the decision verges among continuing with full or reduced computation. The proposed architecture allows to control both the number of blocks that are evaluated, through the use of Early Exits, and the amount of computation conducted within each block. This is conducted by using a clustering algorithm over a discretized space of possibilities.